OpenStudy (anonymous):

d/d+4 + d/d+9 = 1 solve for d . HELP ME .

OpenStudy (anonymous):

d^2+9d+d^2+4d/(d+4)(d+9)=1 2d^2+13d/(d+4)(d+9)=1 2d^2+13d=(d+4)(d+9) 2d^2+13d=d^2+13d+36 d^2=36 sqrt both sides u will get d=6